Two point charges are separated by a distance r. If the separation n is reduced by a factor of 1.5, by what factor does the electric force between them change?
1
Expert's answer
2019-12-10T09:56:57-0500
"F\sim r^{-2}"

Thus,

"\frac{F'}{F}=\frac{r^2}{r'^2}=1.5^2=2.25"